# Teleportation = Translation: Continuous recovery of black hole information

### v1

@@ 1 Introduction
In this work, we construct a smooth unitary interpolation to connect the discrete canonical shift with continuous spacetime evolution. To handle the Type III structure, we employ the Haagerup-Kosaki framework Connes80 ; Kosaki84 and lift the inclusion to a Type II crossed product via Takesaki duality Connes80 ; TakesakiII . Since a trace-preserving conditional expectation is well-defined in this setting, we can construct a strongly continuous unitary path connecting the identity to the canonical shift. We define the continuous generator $\widetilde{G}$ as the unique self-adjoint operator generating this flow.
Building on this construction, we derive our main result by proving that this generator is physically identified with the modular momentum. Specifically, we establish a rigorous identity asserting that the generator of the teleportation flow equals twice the difference of the modular Hamiltonians. The factor of 2 is dictated by Borchers’ commutation relations Borchers92 , ensuring consistency with the scaling of modular flows. Our proof establishes this equality on a dense core of analytic modular vectors and demonstrates the equality of the closed operators via Nelson’s analytic vector theorem Nelson59 . Furthermore, we quantify the stability of this identification under small deformations using noncommutative $L^{p}$-theory Kosaki84 ; Kato . Physically, this result provides a precise operator-algebraic realization of the Teleportation=Translation proposal, demonstrating that information recovery manifests as a continuous geometric translation. Finally, to facilitate the verification of this operator identity in complex physical models where direct analysis is challenging, we propose a concrete observable test based on correlation functions.
The remainder of this paper is organized as follows. In Sec. 2, we summarize the necessary elements of Tomita–Takesaki theory and introduce the Jones projection together with the canonical shift $\Gamma$. In Sec. 3, we construct the continuous unitary interpolation by lifting the Type III inclusion to the crossed product via the Haagerup-Kosaki framework. Sec. 4 is devoted to the proof of our main result, identifying the generator of this flow with the modular momentum ($\widetilde{G}=2P$). We conclude with a discussion of the results in Sec. 5.
@@ 2 The Discrete Protocol: Canonical Shift
2.1 Preliminaries: Tomita-Takesaki Theory
We work on a fixed Gelfand-Naimark-Segal (GNS) Hilbert space $\mathcal{H}$ built from a faithful normal state $\omega$ (with cyclic and separating vector $|\Omega\rangle$) for the von Neumann algebra $\mathcal{M}$ TakesakiI . Let $\mathcal{N}\subset\mathcal{M}$ be an inclusion of von Neumann algebras. For an algebra $\mathcal{X}$ (where $\mathcal{X}=\mathcal{M}$ or $\mathcal{N}$), the core objects of Tomita-Takesaki theory TakesakiII are:
@@ 2.2 The Canonical Shift Protocol
Information is not lost, but relocated from one relative commutant to another.
@@ 3.2 Path A: Pedagogical (CP-Map) Interpolation
3.3 Path B: The Canonical Haagerup-Kosaki Interpolation
In physical settings such as algebraic quantum field theory, the relevant operator algebras–typically denoted $\mathcal{M}$ and $\mathcal{N}$–are of Type III. A hallmark of such algebras is the absence of a normal, faithful, tracial state. Physically, this reflects the fact that the vacuum state exhibits unbounded entanglement across spatial regions, rendering it impossible to assign finite information measures in the usual way. Mathematically, this precludes the existence of a trace-preserving conditional expectation $E:\mathcal{M}\rightarrow\mathcal{N}$, which would otherwise be a natural candidate for an interpolation path. Instead, one must rely on a weaker structure: a faithful normal operator-valued weight (OVW) $\mathcal{E}:\mathcal{M}^{+}\rightarrow\mathcal{N}^{+}$, which lacks the boundedness and trace properties required for our purposes.
To resolve this structural deficit, we employ the continuous crossed-product construction, a cornerstone of the Haagerup-Kosaki framework. This procedure embeds the original Type III algebra $\mathcal{M}$ into a larger semifinite algebra $\tilde{\mathcal{M}}$, typically of Type $\mathrm{II}_{\infty}$, via Takesaki duality. Crucially, this dual algebra admits a canonical semifinite trace $\tau$. Within this lifted setting, the OVW $\mathcal{E}$ is regularized into a genuine conditional expectation $\tilde{E}:\tilde{\mathcal{M}}\rightarrow\tilde{\mathcal{N}}$, which satisfies both the trace-preserving property and idempotency. Conceptually, this lifting introduces an auxiliary structure—often interpreted as a collective coordinate or emergent degree of freedom—that enables us to define a meaningful interpolation path in the absence of a trace in the physical system. Such techniques have recently been leveraged in holographic contexts, including AdS/CFT and black hole entropy dynamics Witten22 .
@@ 3.3 Path B: The Canonical Haagerup-Kosaki Interpolation
The Haagerup-Kosaki theory TakesakiII ; Kosaki84 implements this by constructing the crossed-product algebra using a faithful normal semifinite weight $\omega$ on $\mathcal{M}$ that is chosen to be compatible with the inclusion (i.e., $\omega=\psi\circ\mathcal{E}$ for some weight $\psi$ on $\mathcal{N}$). Note that the existence of such a weight is guaranteed by Haagerup’s theory, and this choice does not result in any loss of generality regarding the structure of the crossed product. Using its associated modular automorphism group $\sigma^{\omega}$, which leaves $\mathcal{N}$ invariant, we construct the crossed-product algebras:
@@
Let $\phi_{0}$ be the faithful normal semifinite weight on the crossed product algebra $\tilde{\mathcal{M}}$ (specifically, the dual weight of the physical state $\omega$). (Note: $\phi_{0}$ is distinct from the canonical trace $\tau$; while $\tau$ provides the structural background for $\tilde{E}$, $\phi_{0}$ carries the physical information). We define the endpoint weight $\phi_{1}:=\phi_{0}\circ\tilde{E}$, which corresponds to the weight associated with the subalgebra $\tilde{\mathcal{N}}$. For $s\in[0,1]$, the canonical $L^{p}$-interpolation constructs a family of conditional expectations $\tilde{E}_{s}:\tilde{\mathcal{M}}\to\tilde{\mathcal{M}}$. We define the one-parameter family of algebras $\tilde{\mathcal{N}}(s)$ as the image of these expectations:
@@ 3.4 The Unitary Path and its Generator
  • Pedagogical Path: The naive CP-interpolation is not idempotent. One cannot define a global Tomita-Takesaki structure. At best, one can analyze correlator-level diagnostics (as in Sec. 4.4) under the assumption that its $s=0$ generator matches the physical one.
  • Physical Path: The semifinite-envelope path $\tilde{E}_{s}$ is idempotent (Theorem 3.6). Its image $\tilde{\mathcal{N}}(s)$ is a true von Neumann subalgebra. Hence, the Tomita-Takesaki machinery applies globally at each $s$. (For a simple demonstration of the unitary goal in a Type I model, see Appendix B.)
@@
The path $\tilde{U}(s)$ does not generally satisfy the group property $\tilde{U}(s+t)=\tilde{U}(s)\tilde{U}(t)$. Therefore, we cannot directly invoke Stone’s Theorem to define a generator $G$ such that $\tilde{U}(s)=e^{-isG}$. Instead, we rigorously define $\tilde{G}$ as the infinitesimal generator (tangent vector) at $s=0$. However, for this definition to be meaningful, we must ensure that the path is sufficiently smooth (differentiable) and that the resulting operator is self-adjoint.
@@
The proof relies on the analytic structure of the Haagerup-Kosaki interpolation and the properties of the modular core.
@@ 4.2 Rigorous Derivation via Modular Perturbation Theory
The result $K^{\prime}(0)=-P$ implies that the infinitesimal change of the modular Hamiltonian with respect to $s$ is exactly generated by the spacetime translation operator $P$. Physically, as we vary $s$, the evolution of the operator $K(s)$ is driven by a real geometric translation in spacetime. This confirms that our continuous interpolation path is aligned with an actual spacetime flow: at $s=0$, the direction in which the modular Hamiltonian initially changes is precisely along the generator $P$ (which translates information outward in the emergent space). Thus, the abstract path $\tilde{E}_{s}$ we construct is not just algebraically idempotent, but its tangent direction is that of a physical translation, substantiating the interpretation of the channel as a true spacetime motion of information.
@@ 4.3 Stability Quantification via Non-Commutative $L^{p}$ Spaces
A crucial physical question is whether the identity $\tilde{G}=2P$ is merely an artifact of our idealized mathematical construction at $s=0$, or if it is structurally robust. Specifically, we must ensure that the generator does not fluctuate wildly under small perturbations of the interpolation parameter $s$. We provide a definitive answer using Kosaki’s non-commutative $L^{p}$ theory Kosaki84 ; Kato , which offers the canonical geometric framework for analyzing perturbations in Type $\text{II}_{\infty}$ von Neumann algebras.
@@ 4.4 Correlation-Function Test of the Conjecture
The equality of the two expressions suggests that the algebraic dynamics of teleportation and the geometric dynamics of spacetime translation are two manifestations of the same underlying phenomenon. In the bulk geometry of semiclassical gravity, the generator $P=K_{\mathcal{M}}-K_{\mathcal{N}}$ acts as the operator conjugate to the time shift across the horizon (specifically, the relative modular Hamiltonian). Physically, perturbations generated by $P$ correspond to injecting energy near the horizon, creating a gravitational shock wave that modifies the background geometry. This shock wave induces a physical time delay (or Shapiro delay) for any incoming matter, which is precisely the mechanism governing the correlation function Shenker2014 ; Gao2017 .
@@ Appendix B Type I Model: Path Failure and Unitary Interpolation Goal
Since $E_{0.5}^{2}(A)\neq E_{0.5}(A)$, this demonstrates the fatal failure of static idempotency for the naive path. This failure is precisely why the Haagerup-Kosaki lift is required in the Type III setting.
@@
Appendix C Haagerup-Kosaki Lift: Existence and Mechanism of $\tilde{E}$
The most critical technical challenge in defining the idempotent path $\tilde{E}_{s}$ in the Type III setting is the construction of a faithful normal conditional expectation. In general, Type III inclusions do not admit a conditional expectation $\mathcal{M}\to\mathcal{N}$. The theory of Haagerup and Takesaki provides the necessary mechanism by lifting the problem from a non-tracial Type III algebra (equipped with an Operator-Valued Weight, OVW) to a tracial semifinite algebra where a genuine conditional expectation $\tilde{E}$ exists. This lift effectively regularizes the unbounded OVW $\mathcal{E}$ via the crossed product construction.
@@ Appendix C Haagerup-Kosaki Lift: Existence and Mechanism of $\tilde{E}$
Let $\mathcal{N}\subset\mathcal{M}$ be an inclusion of von Neumann algebras. While a conditional expectation may not exist, Haagerup’s theorem guarantees the existence of a faithful normal operator-valued weight (OVW) $\mathcal{E}:\mathcal{M}^{+}\to\mathcal{N}^{+}$.
@@
    The crossed product algebra $\tilde{\mathcal{M}}=\mathcal{M}\rtimes_{\sigma^{\omega}}\mathbb{R}$ is a semifinite von Neumann algebra and admits a canonical faithful normal semifinite trace $\tau$ satisfying $\tau\circ\theta_{t}=e^{-t}\tau$, where $\theta_{t}$ is the dual action.
@@
    There exists a unique, faithful normal, semifinite conditional expectation
@@
    Takesaki Duality and Trace: Takesaki duality establishes that $\tilde{\mathcal{M}}$ is a semifinite algebra equipped with a canonical trace $\tau$. This trace allows for the existence of conditional expectations that are not available in the Type III setting.

### v2

@@ 1 Introduction
In this work, we construct a smooth unitary interpolation to connect the discrete canonical shift with continuous spacetime evolution. To handle the Type III structure, we employ the Haagerup–Kosaki framework Connes80 ; Kosaki84 and lift the inclusion to a Type II crossed product via Takesaki duality Connes80 ; TakesakiII . Since a trace-preserving conditional expectation is well-defined in this setting, we can construct a strongly continuous unitary path connecting the identity to the canonical shift. We define the continuous generator $\widetilde{G}$ as the unique self-adjoint operator associated with this flow.
Building on this construction, we derive our main result by proving that this generator is physically identified with the modular momentum. Specifically, we establish a rigorous identity asserting that the generator of the teleportation flow equals twice the difference of the modular Hamiltonians. The factor of 2 is dictated by Borchers’ commutation relations Borchers92 , ensuring consistency with the scaling of modular flows. Our proof establishes this equality on a dense core of analytic modular vectors and demonstrates the equality of the closed operators via Nelson’s analytic vector theorem Nelson59 . Furthermore, we quantify the stability of this identification under small deformations using non-commutative $L^{p}$-theory Kosaki84 ; Kato . Physically, this result provides a precise operator-algebraic realization of the Teleportation=Translation proposal, demonstrating that information recovery manifests as a continuous geometric translation. Finally, to facilitate the verification of this operator identity in complex physical models where direct analysis is challenging, we propose a concrete observable test based on correlation functions.
The remainder of this paper is organized as follows. In Sec. 2, we summarize the necessary elements of Tomita–Takesaki theory and introduce the Jones projection together with the canonical shift $\Gamma$. In Sec. 3, we construct the continuous unitary interpolation by lifting the Type III inclusion to the crossed product via the Haagerup–Kosaki framework. Sec. 4 is devoted to the proof of our main result, identifying the generator of this flow with the modular momentum ($\widetilde{G}=2P$). We conclude with a discussion of the results in Sec. 5.
@@ 2 The Discrete Protocol: Canonical Shift
2.1 Preliminaries: Tomita–Takesaki Theory
We work on a fixed Gelfand-Naimark-Segal (GNS) Hilbert space $\mathcal{H}$ built from a faithful normal state $\omega$ (with cyclic and separating vector $|\Omega\rangle$) for the von Neumann algebra $\mathcal{M}$ TakesakiI . Let $\mathcal{N}\subset\mathcal{M}$ be an inclusion of von Neumann algebras. For an algebra $\mathcal{X}$ (where $\mathcal{X}=\mathcal{M}$ or $\mathcal{N}$), the core objects of Tomita–Takesaki theory TakesakiII are:
@@ 2.2 The Canonical Shift Protocol
Information is not lost, but relocated from one relative commutant to another. Crucially, while explicitly defined algebraically, this shift operator $U_{\Gamma}$ corresponds to a geometric translation in the emergent spacetime—a connection that will be rigorously formulated in the continuum limit in the subsequent sections.
@@ 3.2 Path A: Pedagogical (CP-Map) Interpolation
3.3 Path B: The Canonical Haagerup–Kosaki Interpolation
In physical settings such as algebraic quantum field theory, the relevant operator algebras–typically denoted $\mathcal{M}$ and $\mathcal{N}$–are of Type III. A hallmark of such algebras is the absence of a normal, faithful, tracial state. Physically, this reflects the fact that the vacuum state exhibits unbounded entanglement across spatial regions, rendering it impossible to assign finite information measures in the usual way. Mathematically, this precludes the existence of a trace-preserving conditional expectation $E:\mathcal{M}\rightarrow\mathcal{N}$, which would otherwise be a natural candidate for an interpolation path. Instead, one must rely on a weaker structure: a faithful normal operator-valued weight (OVW) $\mathcal{E}:\mathcal{M}^{+}\rightarrow\mathcal{N}^{+}$, which lacks the boundedness and trace properties required for our purposes. Here, $\mathcal{M}^{+}$ and $\mathcal{N}^{+}$ are positive cones of $\mathcal{M}$ and $\mathcal{N}$, respectively.
To resolve this structural deficit, we employ the continuous crossed-product construction, a cornerstone of the Haagerup–Kosaki framework. This procedure embeds the original Type III algebra $\mathcal{M}$ into a larger semifinite algebra $\tilde{\mathcal{M}}$, typically of Type $\mathrm{II}_{\infty}$, via Takesaki duality. Crucially, this dual algebra admits a canonical semifinite trace $\tau$. Within this lifted setting, the OVW $\mathcal{E}$ is regularized into a genuine conditional expectation $\tilde{E}:\tilde{\mathcal{M}}\rightarrow\tilde{\mathcal{N}}$, which satisfies both the trace-preserving property and idempotency. Conceptually, this lifting introduces an auxiliary structure—often interpreted as a collective coordinate or emergent degree of freedom—that enables us to define a meaningful interpolation path in the absence of a trace in the physical system. Such techniques have recently been leveraged in holographic contexts, including AdS/CFT and black hole entropy dynamics Witten22 .
@@ 3.3 Path B: The Canonical Haagerup–Kosaki Interpolation
The Haagerup–Kosaki theory TakesakiII ; Kosaki84 implements this by constructing the crossed-product algebra using a faithful normal semifinite weight $\omega$ on $\mathcal{M}$ that is chosen to be compatible with the inclusion (i.e., $\omega=\psi\circ\mathcal{E}$ for some weight $\psi$ on $\mathcal{N}$). Note that the existence of such a weight is guaranteed by Haagerup’s theory, and this choice does not result in any loss of generality regarding the structure of the crossed product. Using its associated modular automorphism group $\sigma^{\omega}$, which leaves $\mathcal{N}$ invariant, we construct the crossed-product algebras:
@@
Let $\phi_{0}$ be the faithful normal semifinite weight on the crossed product algebra $\tilde{\mathcal{M}}$ (specifically, the dual weight of the physical state $\omega$). (Note: $\phi_{0}$ is distinct from the canonical trace $\tau$; while $\tau$ provides the structural background for $\tilde{E}$, $\phi_{0}$ carries the physical information). We define the endpoint weight $\phi_{1}:=\phi_{0}\circ\tilde{E}$, which corresponds to the weight associated with the subalgebra $\tilde{\mathcal{N}}$. This relation formally mirrors the compatibility condition $\omega=\psi\circ\mathcal{E}$ at the level of dual weights. Crucially, however, whereas $\omega$ was employed to guarantee the inclusion property during the lifting process, $\phi_{1}$ serves as the coarse-grained target weight on $\tilde{\mathcal{N}}$ that defines the interpolation trajectory.
For $s\in[0,1]$, the canonical $L^{p}$-interpolation constructs a family of conditional expectations $\tilde{E}_{s}:\tilde{\mathcal{M}}\to\tilde{\mathcal{M}}$. We define the one-parameter family of algebras $\tilde{\mathcal{N}}(s)$ as the image of these expectations:
@@ 3.4 The Unitary Path and its Generator
  • Pedagogical Path: The naive CP-interpolation is not idempotent. One cannot define a global Tomita–Takesaki structure. At best, one can analyze correlator-level diagnostics (as in Sec. 4.4) under the assumption that its $s=0$ generator matches the physical one.
  • Physical Path: The semifinite-envelope path $\tilde{E}_{s}$ is idempotent (Theorem 3.6). Its image $\tilde{\mathcal{N}}(s)$ is a true von Neumann subalgebra. Hence, the Tomita–Takesaki machinery applies globally at each $s$. (For a simple demonstration of the unitary goal in a Type I model, see Appendix B.)
@@
The path $\tilde{U}(s)$ does not generally satisfy the group property $\tilde{U}(s+t)=\tilde{U}(s)\tilde{U}(t)$. Therefore, Stone’s Theorem does not directly apply to define a generator $\tilde{G}$ such that $\tilde{U}(s)=e^{-is\tilde{G}}$. Instead, we rigorously define $\tilde{G}$ as the infinitesimal generator (tangent vector) at $s=0$. However, for this definition to be meaningful, we must ensure that the path is sufficiently smooth (differentiable) and that the resulting operator is self-adjoint. Let $\mathcal{D}_{\text{core}}$ be the dense set of vectors that are analytic for the modular flow $\sigma_{t}$.
@@
The immediate physical consequence of this core preservation is that the constructed path is sufficiently smooth to define a generator. Specifically, the invariant core ensures regularity at the modular boundary:
@@
The proof relies on the analytic structure of the Haagerup–Kosaki interpolation and the properties of the modular core.
@@ 4.2 Rigorous Derivation via Modular Perturbation Theory
The result $K^{\prime}(0)=-P$ implies that the infinitesimal change of the modular Hamiltonian with respect to $s$ is exactly generated by the spacetime translation operator $P$. As we vary $s$, the evolution of the operator $K(s)$ is driven by a real geometric translation in spacetime. This confirms that our continuous interpolation path is aligned with an actual spacetime flow: at $s=0$, the direction in which the modular Hamiltonian initially changes is precisely along the generator $P$ (which translates information outward in the emergent space). Thus, the abstract path $\tilde{E}_{s}$ we construct is not just algebraically idempotent, but its tangent direction is that of a physical translation, substantiating the interpretation of the channel as a true spacetime motion of information.
@@ 4.3 Stability Quantification via Non-Commutative $L^{p}$ Spaces
A crucial physical question is whether the identity $\tilde{G}=2P$ is merely an artifact of our idealized mathematical construction at $s=0$, or if it is structurally robust. Specifically, we must ensure that the generator does not fluctuate wildly under small perturbations of the interpolation parameter $s$. We provide a definitive answer using Kosaki’s non-commutative $L^{p}$ theory Kosaki84 ; Kato , which offers the canonical geometric framework for analyzing perturbations in Type II von Neumann algebras.
@@ 4.4 Correlation-Function Test of the Conjecture
The equality of the two expressions suggests that the algebraic dynamics of teleportation and the geometric dynamics of spacetime translation are two manifestations of the same underlying phenomenon. In the bulk geometry of semiclassical gravity, the generator $P=K_{\mathcal{M}}-K_{\mathcal{N}}$ acts as the operator conjugate to the time shift across the horizon (specifically, the relative modular Hamiltonian). Perturbations generated by $P$ correspond to injecting energy near the horizon, creating a gravitational shock wave that modifies the background geometry. This shock wave induces a physical time delay (or Shapiro delay) for any incoming matter, which is precisely the mechanism governing the correlation function Shenker2014 ; Gao2017 .
@@ Appendix B Type I Model: Path Failure and Unitary Interpolation Goal
Since $E_{0.5}^{2}(A)\neq E_{0.5}(A)$, this demonstrates the fatal failure of static idempotency for the naive path. This failure is precisely why the Haagerup–Kosaki lift is required in the Type III setting.
@@
Appendix C Haagerup–Kosaki Lift: Existence and Mechanism of $\tilde{E}$
The most critical technical challenge in defining the idempotent path $\tilde{E}_{s}$ in the Type III setting is the construction of a faithful normal conditional expectation. In general, Type III inclusions do not admit a conditional expectation $\mathcal{M}\to\mathcal{N}$. The theory of Haagerup and Takesaki provides the necessary mechanism by lifting the problem from a non-tracial Type III algebra (equipped with an OVW) to a tracial semifinite algebra where a genuine conditional expectation $\tilde{E}$ exists. This lift effectively regularizes the unbounded OVW $\mathcal{E}$ via the crossed product construction.
@@ Appendix C Haagerup–Kosaki Lift: Existence and Mechanism of $\tilde{E}$
Let $\mathcal{N}\subset\mathcal{M}$ be an inclusion of von Neumann algebras. While a conditional expectation may not exist, Haagerup’s theorem guarantees the existence of a faithful normal OVW $\mathcal{E}:\mathcal{M}^{+}\to\mathcal{N}^{+}$.
@@
[…]
    $\tau\circ\theta_{t}=e^{-t}\tau.$ (38)
@@
    There exists a unique, faithful normal conditional expectation
@@
    Takesaki Duality and Trace: Takesaki duality establishes that $\tilde{\mathcal{M}}$ is a semifinite algebra equipped with a canonical trace $\tau$. The scaling property (38) is characteristic of the Type II crossed product derived from a Type III algebra, allowing us to treat modular dynamics as trace-scaling automorphisms. Consequently, the existence of this trace allows for the existence of conditional expectations that are not available in the Type III setting.